# Aa Article Title

### v1

@@ 2.1 Splitting Criterion
for inputs $\bm{X}_{i}\in\mathbb{R}^{d_{1}\times d_{2}}$ and responses $y_{i}\in\mathbb{R}$. Here $g^{*}\colon\mathbb{R}^{d_{1}\times d_{2}}\rightarrow\mathbb{R}$ is a real-valued function we wish to estimate and the $E_{i}$s are independent mean zero noises. Extending the tree model (2) and creating ensembles for tensor input regression require ingredients from low-rank tensor approximations (i.e., (7) and (8)) and associated low-rank tensor input regression models.
@@
    $R_{1}\coloneqq\left\{\bm{X}\mid\bm{X}[:,j_{1},j_{2}]\leq\bm{X}[j_{0},j_{1},j_{2}]\right\},$ (left child) (10)
    $R_{2}\coloneqq\left\{\bm{X}\mid\bm{X}[:,j_{1},j_{2}]>\bm{X}[j_{0},j_{1},j_{2}]\right\},$ (right child) (11)
    minimizes the sum of variances
    $\text{SSE}(j_{0},j_{1},j_{2})=\sum_{j=1}^{2}\frac{1}{N_{j}}\sum_{\bm{X}_{i}\in R_{j}}(y_{i}-\hat{y}_{R_{j}})^{2}$ $=\sum_{j=1}^{2}\frac{1}{N_{j}}\sum_{\underset{I_{j}=\{k\mid\bm{X}[k,j_{1},j_{2}]\in R_{j}\}}{i\in I_{j}}}\left\|\bm{y}[I_{j},:]-\frac{1}{|I_{j}|}\bm{1}_{|I_{j}|}^{\top}\bm{y}[I_{j},:]\cdot\bm{1}_{|I_{j}|}\right\|_{2}^{2}$ (12)
    among all possible pairs of children induced by splitting on a dimension pair and observed value.
  2. 2.
    Low-rank approximation error (LAE). This criterion aims to partition the input space by leveraging the potential low-rank structure in the tensor input $\bm{X}$. Rather than use the predictors, we can split based on how closely the children tensor inputs match their low-rank CP or Tucker approximations. Namely, we choose the child pair (10) and (11) whose tensor inputs $\bm{X}[I_{j},:,:]\in\mathbb{R}^{|I_{j}|\times d_{1}\times d_{2}}$ (for $j=1,2$) and corresponding tensor low-rank approximations $\widetilde{\bm{X}}[I_{j},:,:]$ minimize the error
    $\text{LAE}(j_{0},j_{1},j_{2})=\sum_{j=1}^{2}\sum_{\underset{I_{j}=\{k\mid\bm{X}[k,j_{1},j_{2}]\in R_{j}\}}{i\in I_{j}}}\left\|\widetilde{\bm{X}}[i,:,:]-\bm{X}[i,:,:]\right\|_{F}^{2}$ (13)
@@
    $\text{LRE}(j_{0},j_{1},j_{2})=\sum_{j=1}^{2}\sum_{\underset{I_{j}=\{k\mid\bm{X}[k,j_{1},j_{2}]\in R_{j}\}}{i\in I_{j}}}\left\|\widetilde{\bm{y}}\left(\bm{X}[i,:,:]\right)-y_{i}\right\|_{F}^{2},$ (14)
    where $\widetilde{\bm{y}}\left(\bm{X}[I_{j},:,:]\right)$ is the chosen CP (7) or Tucker (8) low-rank regression model (of a given target rank) of $\bm{X}[I_{j},:,:]$ against $\bm{y}[I_{j}]$. Intuitively, this minimizes the error of predicting $\bm{y}$ in each child, which is analogous to CART minimizing the sum of variances of $\bm{y}$ in each node. Because we fit two models for two children nodes, (14) is slightly more expensive than (13).
Any given split with $D=3$ requires a split-coordinate pair $(j_{1},j_{2})$ and index $j_{0}$ for the split value $c=\bm{X}[j_{0},j_{1},j_{2}]$. The $j_{0},j_{1},j_{2}$ triplet is usually obtained by solving the following mixed integer problem of dimensionality $nd_{1}d_{2}$:
@@
The splitting criteria SSE and LRE involve both $\bm{X}$ and $\bm{y}$, whereas LAE is specifically designed for low-rank tensors $\bm{X}$ (or mat-vec that essentially has low-rank structures). Because LAE does not involve the response $\bm{y}$, the learned tree based on LAE will well approximate the mean response only when the underlying latent tree structure in the tensor input $\bm{X}$ corresponds to that in $\bm{y}$. On the other hand, SSE and LRE more directly target the prediction task and more easily produce a good fit to the data, but also make the algorithm more prone to overfitting, though this issue can be easily addressed with proper regularization techniques such as pruning and out-of-sample validation.
@@
The key idea behind the LRE loss is to incorporate low-rank tensor regression as the baseline models on the leaves of a partition tree. Trees can capture non-linear local structures in the underlying mean functions but are ineffective at approximating global smooth structures such as linear patterns on some latent dimensions which low-rank tensor regression targets. As such, we construct trees using novel splitting criteria and complexity measures, then fit tensor regressions at each leaf node. Figure 1 shows a scenario of mixed class data (Zhou et al., 2013) in which existing tensor regressions (i.e., the top row) fail to recover each class but a tree-based regression model is successful with appropriate splits. In addition, this tree-based model can realize typical low-rank CP and Tucker regression models as special cases when the max depth of the tree is set to be zero (i.e., when the tree has only one node) and the only model is a single low-rank regression model. The effects of the splitting criteria will be further investigated in Section 5.1.
Result: Train a tensor input single decision tree regressor with exhaustive search
Function fit($\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$, $\bm{y}\in\mathbb{R}^{n}$):
begin
       Initialize root node. for node in tree do
             Calculate for all potential splits along each dimension pair $(j_{1},j_{2})$: Find $\arg\min_{j_{0}\in\{1,\cdots,n\},j_{1}\in\{1,\cdots,d_{1}\},j_{2}\in\{1,\cdots,d_{2}\}}LAE(j_{0},j_{1},j_{2})$ or $LRE(j_{0},j_{1},j_{2})$ ;
             With the minimizer $(j_{0}^{*},j_{1}^{*},j_{2}^{*})$, split the dataset on the chosen dimension pair $(j_{1}^{*},j_{2}^{*})$ and split value $\bm{X}[j_{0}^{*},j_{1}^{*},j_{2}^{*}]$, creating left and right child nodes using data in $R_{1},R_{2}$ ;
       end for
      Fit the chosen mean/CP/Tucker models at each of the leaf nodes.
end
Algorithm 1 Exhaustive search method for fitting tensor input decision tree regressor with low-rank splitting criteria (13) and (14).
@@ 2.2 Reducing Computational Cost
There are two major computational overheads in fitting the tree tensor regression model we just proposed: leaf model fitting and splitting criteria computation. The cost of the former is alleviated thanks to the tree structure providing a partitioning regime that allows divide-and-conquer scaling. The latter, however, is a new computational challenge born from the nature of tensor data as we explain below.
It is computationally very expensive to exhaustively search (see Algorithm 1) for the best split node in tensor input decision trees. This process evaluates all possible splits along each dimension pair $(j_{1},j_{2})$, which requires examining every combination of $j_{0}$, $j_{1}$, and $j_{2}$. Consequently, the search space size for each split decision is $n\cdot d_{1}\cdot d_{2}$, which scales poorly with increasing dataset size and tensor dimensions, making exhaustive search computationally intensive and impractical for large-scale data or real-time applications.
@@
Computing the variance in (12) using scalar responses has complexity $\mathcal{O}\left(n\right)$, which can be vectorized and does not scale with dimensions. On the other hand, the time complexities of (13) and (14) are dominated by the CP and Tucker low-rank decomposition of the coefficient tensors. Although many scalable methods are proposed for CP and Tucker decompositions, the well-adopted alternating least square (ALS) (Malik and Becker, 2018) solvers are of iterative nature and hence can have relatively high complexities:
@@
(CP-ALS per iteration complexity Minster et al. (2023) Section 3.2) In the alternating least square algorithm for the rank-$R$ CP decomposition of a $n\times d_{1}\times d_{2}\times\cdots\times d_{K}$ tensor where $R<n$ and $R\ll d_{i},i=1,\cdots,K$, each iteration has time complexity $\mathcal{O}\left((K+1)\cdot\prod_{i=1}^{K}d_{i}\cdot R\right)$. Therefore, for $N^{*}$ iterations in CP-ALS, the overall complexity is bounded from above by $\mathcal{O}\left(N^{*}\cdot(K+1)R\cdot\left(\max_{i}d_{i}\right)^{K}\right)$.
@@
When $R<n$ and $d_{i}^{{}^{\prime}}\asymp d_{i}\ll n,i=1,\cdots,K$, for $N^{*}$ iterations in Tucker-ALS, the overall complexity is at most $\mathcal{O}\left(N^{*}\cdot n\cdot\left(\max_{i}d_{i}\right)^{2K}\right)$.
To simplify our subsequent analysis, we assume $n\asymp R$ and always choose a constant maximal iteration number $N^{*}$ regardless of the convergence. Then both decompositions have time complexities at most $\mathcal{O}\left(N^{*}\cdot n\cdot\left(\max_{i}d_{i}\right)^{2K}\right)$ which scales linearly with $n$. However, the factor $N^{*}\cdot\left(\max_{i}d_{i}\right)^{2K}$ greatly increases complexity in greedy-style tree-fitting, which bottlenecks most regression tree fitting algorithms (Loh, 2014).
@@
Minimizing these three criteria can be a strategy to choose the split-value index $j_{0}$ and splitting coordinates $(j_{1},j_{2})$, and can easily generalize to regression trees for tensor inputs. Finding the best combination of $(j_{0},j_{1},j_{2})$ using (3), (13), (14), in a worst-case scenario, needs $n\cdot\prod_{j=1}^{K}d_{j}$-many variance evaluations and contributes to a quadratic complexity in $n$. But if we do not optimize over $j_{0}$ and instead split on the sample mean, we can simplify the partitions in (3), (13), (14) as:
$\bar{R}_{1}\coloneqq\left\{\bm{X}\mid\bm{X}[:,j_{1},j_{2}]\leq\frac{1}{n}\sum_{k=1}^{n}\bm{X}[k,j_{1},j_{2}]\right\},$ (left child) (16)
$\bar{R}_{2}\coloneqq\left\{\bm{X}\mid\bm{X}[:,j_{1},j_{2}]>\frac{1}{n}\sum_{k=1}^{n}\bm{X}[k,j_{1},j_{2}]\right\}.$ (right child) (17)
Since we no longer need to determine $j_{0}$, we can complete the search within linear time in $n$, without losing too much empirical prediction performance. Then the corresponding loss functions in our reduced optimization problems become:
$\overline{\text{SSE}}(j_{1},j_{2})=\sum_{j=1}^{2}\frac{1}{|I_{j}|}\sum_{\underset{I_{j}=\{k\mid\bm{X}[k,j_{1},j_{2}]\in\bar{R}_{j}\}}{i\in I_{j}}}\left\|\hat{y}_{\bar{R}_{j}}-y_{i}\right\|_{2}^{2},$ (18)
$\overline{\text{LAE}}(j_{1},j_{2})=\sum_{j=1}^{2}\sum_{\underset{I_{j}=\{k\mid\bm{X}[k,j_{1},j_{2}]\in\bar{R}_{j}\}}{i\in I_{j}}}\left\|\widetilde{\bm{X}}[i,:,:]-\bm{X}^{(n)}[i,:,:]\right\|_{F}^{2},$ (19)
$\overline{\text{LRE}}(j_{1},j_{2})=\sum_{j=1}^{2}\sum_{\underset{I_{j}=\{k\mid\bm{X}[k,j_{1},j_{2}]\in\bar{R}_{j}\}}{i\in I_{j}}}\left\|\widetilde{\bm{y}}\left(\bm{X}^{(n)}[i,:,:]\right)-y_{i}\right\|_{F}^{2}.$ (20)
This simplification shrinks the dimensionality of the optimization problem $\min_{(j_{0},j_{1},j_{2})}\mathcal{L}(j_{0},j_{1},j_{2})$ to reduce the complexity per evaluation of the split criterion, by removing the search along the first mode.
Figure 2 shows that the loss functions LAE and LRE lead to similar complexities and predictive behavior. Although LAE and LRE are both based on low-rank tensor decomposition, LRE (like SSE) focuses on the children’s predictive loss. In what follows, we mainly study TT models with LRE in (14) or its mean version (20) as our default loss functions, unless otherwise is stated.
Methods of efficient search for splitting coordinates.
Existing efficient methods (Luo and Pratola, 2022; Chickering et al., 2001) for splitting point optimization (i.e., search among possible points for the above loss functions) usually rely on the distribution assumptions on the candidates. In our case, this assumption is usually not realistic due to the complex correlation between dimensions of the input tensor. To reduce the $N^{*}\cdot\left(\max_{i}d_{i}\right)^{K}$ factor further, we propose the following two methods for shrinking the search space of these optimization problems in order to quickly solve the minimization problems for (13) or (14). Both methods have existed in the optimization community for a while and exhibit trade-offs between efficiency and accuracy, but to our best knowledge are applied in regression trees for the first time.
Item Model Decision Trees Extra-Trees (ERT) Algorithm 3 Algorithm 4 Reference (Breiman et al., 1984) (Geurts et al., 2006) This paper splitting coordinate Exact optimization Exact optimization Importance sampling Branch-and-bound splitting value Exact optimization Random selection Random $\tau<1$/ Random $\xi>0$/ Exact $\tau=1$ Exact $\xi=0$
@@
See Appendix B. ∎
This complexity result indicates that the tensor tree model has $\mathcal{O}(n^{2}\cdot C)$ complexity (where $C=C(d_{1},d_{2})$ depends on the dimensionality of the tensor inputs) compared to the larger $\mathcal{O}(n^{3})$ complexity (whose constant inside big O also depends on the kernel evaluation) incurred by tensor GP models (Yu et al., 2018) as shown later in Figure 8. We can also see that both LAE and LRE share the same big O complexity in terms of the sample size $n$. As explained in Section 2.1, we will focus on LRE since it parallels the SSE in regular regression trees.
Figure 2: Out-of-sample MSEs for tree regression models of different maximal depths, using a $(500,5,4)$ input tensor $\bm{X}$ sampled uniformly randomly from $[-1,1]$ and scalar output $\bm{y}=2\bm{X}[:,0,1]\cdot\bm{X}[:,2,3]+3\bm{X}[:,1,0]\cdot\bm{X}[:,2,0]\cdot\bm{X}[:,3,0]+\varepsilon$ where $\varepsilon\sim N(0,\sigma^{2})$ and $\sigma^{2}=0,0.01,0.1$. (Columns 1&4: Full search; Columns 2&5: LS with $\tau=0.5$; Columns 3&6: BB with $\delta=0.5$.)

### v2

@@ 2 Trees for Tensor Inputs
for inputs $\bm{X}_{i}\in\mathbb{R}^{d_{1}\times d_{2}}$ and responses $y_{i}\in\mathbb{R}$. Here $g^{*}\colon\mathbb{R}^{d_{1}\times d_{2}}\rightarrow\mathbb{R}$ is a real-valued function we wish to estimate and the $E_{i}$s are independent mean zero noises. Extending the tree model (2) and creating ensembles for tensor input regression require ingredients from low-rank tensor approximations (i.e., (5) and (6)) and associated low-rank tensor input regressions.
2.1 Splitting Criterion
@@
$R_{1}\coloneqq\left\{\bm{X}\mid\bm{X}[:,j_{1},j_{2}]\leq\bm{X}[j_{0},j_{1},j_{2}]\right\}\quad R_{2}\coloneqq\left\{\bm{X}\mid\bm{X}[:,j_{1},j_{2}]>\bm{X}[j_{0},j_{1},j_{2}]\right\}$ (8)
minimizes the sum of variances (here $I_{j}\coloneqq\{i\mid\bm{X}[i,j_{1},j_{2}]\in R_{j}\}$ and $\bm{y}[I_{j},:]=\{y_{i}\mid i\in I_{j}\}$)
$\text{SSE}(j_{0},j_{1},j_{2})=\sum_{j=1}^{2}\frac{1}{N_{j}}\sum_{\bm{X}_{i}\in R_{j}}(y_{i}-\hat{y}_{R_{j}})^{2}=\sum_{j=1}^{2}\frac{1}{N_{j}}\left\|\bm{y}[I_{j},:]-\frac{1}{|I_{j}|}\bm{1}_{|I_{j}|}^{\top}\bm{y}[I_{j},:]\cdot\bm{1}_{|I_{j}|}\right\|_{2}^{2}$ (9)
over all such possible pairs of children.
Low-rank approximation error (LAE). This criterion aims to partition the input space by leveraging the potential low-rank structure in the tensor input $\bm{X}$. Rather than use the predictors, we can split based on how closely the children tensor inputs match their low-rank CP or Tucker approximations. Namely, we choose the child pair (8) whose tensor inputs $\bm{X}[I_{j},:,:]\in\mathbb{R}^{|I_{j}|\times d_{1}\times d_{2}}$ (for $j=1,2$) and corresponding tensor low-rank approximations $\widetilde{\bm{X}}[I_{j},:,:]$ minimize the error (again, here $I_{j}\coloneqq\{i\mid\bm{X}[i,j_{1},j_{2}]\in R_{j}\}$)
$\text{LAE}(j_{0},j_{1},j_{2})=\sum_{j=1}^{2}\sum_{i\in I_{j}}\left\|\widetilde{\bm{X}}[i,:,:]-\bm{X}[i,:,:]\right\|_{F}^{2}$ (10)
@@
$\text{LRE}(j_{0},j_{1},j_{2})=\sum_{j=1}^{2}\sum_{i\in I_{j}}\left\|\widetilde{\bm{y}}\left(\bm{X}[i,:,:]\right)-y_{i}\right\|_{F}^{2},$ (11)
where $\widetilde{\bm{y}}\left(\bm{X}[I_{j},:,:]\right)$ is the chosen CP (5) or Tucker (6) low-rank regression model (of a given target rank) of $\bm{X}[I_{j},:,:]$ against $\bm{y}[I_{j}]$. Intuitively, this minimizes the error of predicting $\bm{y}$ in each child, which is analogous to CART minimizing the sum of variances of $\bm{y}$ in each node. Because we fit two models for two children nodes, (11) is slightly more expensive than (10). Any given split with $D=3$ requires a split-coordinate pair $(j_{1},j_{2})$ and index $j_{0}$ for the split value $c=\bm{X}[j_{0},j_{1},j_{2}]$. The $j_{0},j_{1},j_{2}$ triplet is usually obtained by solving the following mixed integer problem of dimensionality $nd_{1}d_{2}$:
@@
The splitting criteria SSE and LRE involve both $\bm{X}$ and $\bm{y}$, whereas LAE is specifically designed for low-rank tensors $\bm{X}$ (or mat-vec that essentially has low-rank structures). Because LAE does not involve the response $\bm{y}$, the learned tree based on LAE will well approximate the mean response only when the underlying latent tree structure in $\bm{X}$ corresponds to that in $\bm{y}$. On the other hand, SSE and LRE more directly target the prediction task and more easily produce a good fit to the data, but also make the algorithm more prone to overfitting, though this issue can be easily addressed with proper regularization techniques such as pruning and out-of-sample validation.
@@
The key idea behind the LRE loss is to incorporate low-rank tensor regression as the baseline models on the leaves of a partition tree. Trees can capture non-linear local structures in the underlying mean functions but are ineffective at approximating global smooth structures such as linear patterns on some latent dimensions which low-rank tensor regression targets. As such, we construct trees using novel splitting criteria and complexity measures, then fit tensor regressions at each leaf node. To show the effectiveness of our model, we first generate three classes of coefficient tensors $\bm{B}_{1},\bm{B}_{2},\bm{B}_{3}$, and then simulate i.i.d. Gaussian tensors $\bm{X}_{i}$ to yield the response $\bm{y}_{i}=\bm{X}_{i}\circ\bm{B}_{i}+\delta_{i}$ for each class, where the elements of $\delta_{i}$ are independent shifted noises $N(i,\sigma^{2})$ to ensure identifiability between these three classes. Figure 1 shows a scenario of mixed class data (Zhou et al., 2013) in which existing tensor regressions (i.e., the top row) fail to recover each class but a tree-based regression model is successful with appropriate splits. In addition, this tree-based model can realize typical low-rank CP and Tucker regression models as special cases when the max depth of the tree is set to be zero (i.e., when the tree has only one node) and the only model is a single low-rank regression model. The effects of the splitting and pruning criteria will be further investigated in Section 5.1 and Appendix D.
As a final note, the variance and LRE splitting criteria can be used to tackle a regression problem with a tensor response and vector predictor. On the other hand, a CP or Tucker decomposition of a vector predictor is just the predictor itself, so the LAE error would be zero for any pair of potential children and hence cannot be used with a vector predictor.
1
Result: Train a tensor-input single decision tree regressor with exhaustive search
2
3Function fit($\bm{X}\in\mathbb{R}^{n\times d_{1}\times d_{2}}$, $\bm{y}\in\mathbb{R}^{n}$):
4begin
5       Initialize root node. for node in tree do
6             Calculate for all potential splits along each dimension pair $(j_{1},j_{2})$: Find $\arg\min_{j_{0}\in\{1,\cdots,n\},j_{1}\in\{1,\cdots,d_{1}\},j_{2}\in\{1,\cdots,d_{2}\}}LAE(j_{0},j_{1},j_{2})$ or $LRE(j_{0},j_{1},j_{2})$ ;
7             With the minimizer $(j_{0}^{*},j_{1}^{*},j_{2}^{*})$, split the dataset on the chosen dimension pair $(j_{1}^{*},j_{2}^{*})$ and split value $\bm{X}[j_{0}^{*},j_{1}^{*},j_{2}^{*}]$, creating left and right child nodes using data in $R_{1},R_{2}$ ;
8
9       end for
10      Fit the chosen mean/CP/Tucker models at each of the leaf nodes.
11 end
12
Algorithm 1 Exhaustive search method for fitting tensor-input decision tree regressor with low-rank splitting criteria (10) and (11).
@@ 2.2 Reducing Computational Cost
There are two major computational overheads in fitting the tree tensor regression model we just proposed: leaf model fitting and splitting criteria computation. The cost of the former is alleviated by the tree structure providing a partitioning regime that allows divide-and-conquer scaling. The latter, however, is a new computational challenge born from the nature of tensor data as we explain below.
It is computationally very expensive to exhaustively search (see Algorithm 1) for the best split node in tensor-input decision trees. This process evaluates all possible splits along each dimension pair $(j_{1},j_{2})$, which requires examining every combination of $j_{0}$, $j_{1}$, and $j_{2}$. Consequently, the search space size for each split decision is $n\cdot d_{1}\cdot d_{2}$, which scales poorly with increasing dataset size and tensor dimensions, making exhaustive search computationally intensive and impractical for large-scale data or real-time applications.
@@
Computing the variance in (9) for scalar responses has complexity $\mathcal{O}\left(n\right)$, which can be vectorized and does not scale with dimensions. On the other hand, the complexities of (10) and (11) are dominated by the CP and Tucker decomposition of the coefficient tensors. Although many scalable methods are proposed for CP and Tucker decompositions, the well-adopted alternating least square (ALS) (Malik and Becker, 2018) solvers are of iterative nature and hence can have relatively high complexities:
@@
(CP-ALS per iteration complexity Minster et al. (2023) Section 3.2) In the alternating least square algorithm for the rank-$R$ CP decomposition of a $n\times d_{1}\times d_{2}\times\cdots\times d_{K}$ tensor where $R<n$ and $R\ll d_{i},i=1,\cdots,K$, each iteration has time complexity $\mathcal{O}((K+1)\cdot\prod_{i=1}^{K}d_{i}\cdot R)$. Therefore, for $N^{*}$ iterations in CP-ALS, the overall complexity is bounded from above by $\mathcal{O}\left(N^{*}\cdot(K+1)R\cdot(\max_{i}d_{i}\right)^{K})$.
@@
When $R<n$ and $d_{i}^{{}^{\prime}}\asymp d_{i}\ll n,i=1,\cdots,K$, for $N^{*}$ iterations in Tucker-ALS, the overall complexity is at most $\mathcal{O}\left(N^{*}\cdot n\cdot(\max_{i}d_{i}\right)^{2K})$.
To simplify our subsequent analysis, we assume $n\asymp R$ and always choose a constant maximal iteration number $N^{*}$ regardless of the convergence. Then both decompositions have time complexities at most $\mathcal{O}(N^{*}\cdot n\cdot\left(\max_{i}d_{i}\right)^{2K})$ which scales linearly with $n$. However, the factor $N^{*}\cdot\left(\max_{i}d_{i}\right)^{2K}$ greatly increases complexity in greedy-style tree-fitting.
@@
Minimizing these three criteria can be a strategy to choose the split-value index $j_{0}$ and splitting coordinates $(j_{1},j_{2})$, and can easily generalize to regression trees for tensor inputs. Finding the best combination of $(j_{0},j_{1},j_{2})$ using (3), (10), (11), in a worst-case scenario, needs $n\cdot\prod_{j=1}^{K}d_{j}$-many variance evaluations and contributes to a quadratic complexity in $n$. But if we do not optimize over $j_{0}$ and instead split on the sample mean $m_{j_{1},j_{2}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\bm{X}[i,j_{1},j_{2}]$, we can simplify the partitions in (3), (10), (11) as:
$\bar{R}_{1}\coloneqq\left\{\bm{X}\mid\bm{X}[:,j_{1},j_{2}]\leq m_{j_{1},j_{2}}\right\},\quad\bar{R}_{2}\coloneqq\left\{\bm{X}\mid\bm{X}[:,j_{1},j_{2}]>m_{j_{1},j_{2}}\right\}.$ (13)
Since we no longer need to determine $j_{0}$, we can complete the search within linear time in $n$, without losing too much empirical prediction performance. Then the corresponding loss functions in our reduced optimization problems become (here $I_{j}\coloneqq\{i\mid\bm{X}[i,j_{1},j_{2}]\in\bar{R}_{j}\}$):
$\overline{\text{SSE}}(j_{1},j_{2})=\sum_{j=1}^{2}\frac{1}{|I_{j}|}\sum_{i\in I_{j}}\left\|\hat{y}_{\bar{R}_{j}}-y_{i}\right\|_{2}^{2},$ (14)
$\overline{\text{LAE}}(j_{1},j_{2})=\sum_{j=1}^{2}\sum_{i\in I_{j}}\left\|\widetilde{\bm{X}}[i,:,:]-\bm{X}^{(n)}[i,:,:]\right\|_{F}^{2},$ (15)
$\overline{\text{LRE}}(j_{1},j_{2})=\sum_{j=1}^{2}\sum_{i\in I_{j}}\left\|\widetilde{\bm{y}}\left(\bm{X}^{(n)}[i,:,:]\right)-y_{i}\right\|_{F}^{2}.$ (16)
Removing the search along the first mode shrinks the dimensionality of the optimization problem $\min_{(j_{0},j_{1},j_{2})}\mathcal{L}(j_{0},j_{1},j_{2})$ to reduce the complexity per evaluation of the split criterion.
Figure 2 shows that the loss functions LAE and LRE lead to similar complexities and predictive behavior. Although both losses are based on low-rank tensor decomposition, LRE (like SSE) focuses on the children’s predictive loss. In what follows, we mainly study TT models with LRE in (11) or its mean version (16) as our default loss functions, unless otherwise stated. Appendix B details two advanced searching techniques, namely leverage score sampling (LS) and branch-and-bound (BB) methods for even more efficient searches.
@@
See Appendix F. ∎
This complexity result indicates that the tensor tree model has $\mathcal{O}(n^{2}\cdot C)$ complexity (where $C=C(d_{1},d_{2})$ depends on the dimensionality of the tensor inputs) compared to the larger $\mathcal{O}(n^{3})$ complexity (whose constant inside big O also depends on the kernel evaluation) incurred by tensor GP models (Yu et al., 2018) as shown later in Figure 6. We can also see that both LAE and LRE share the same big O complexity in terms of $n$. As explained in Section 2.1, we will focus on LRE since it parallels the SSE in regular regression trees.
Figure 2: Out-of-sample MSEs for tree regression models of different maximal depths, using a $(500,5,4)$ input tensor $\bm{X}$ sampled uniformly randomly from $[-1,1]$ and scalar output $\bm{y}=2\bm{X}[:,0,1]\cdot\bm{X}[:,2,3]+3\bm{X}[:,1,0]\cdot\bm{X}[:,2,0]\cdot\bm{X}[:,3,0]+\varepsilon$. Columns 1&4: Full search; Columns 2&5: LS with $\tau=0.5$; Columns 3&6: BB with $\delta=0.5$.